(a) Let X be a random variable with the following probability distribution x -3 6 9 f(x) 1 6 1 2 1 3 Find the mean and variance of Y = 3X + 2. (b) If X ∼ Bin(6, 1 3 ) and Y ∼ N( 1 2 , 1 4 ) are independent random variables, find the mean and variance of the random variable W = 2X − 4Y + 5.


1
Expert's answer
2021-02-24T07:45:04-0500

Solution

a).

"E(y) = 3E(x)+2""E(x) =(3*{1 \over 6}) + (6*{1 \over 2})+(9*{1 \over 3})"

"=6.5"


"E(y) = 3(6.5)+2"

Mean =21.5



"Var(y) = 3^2 var(x) = 9var(x)"

"var(x) = \sum x^2 P(X=x) - {(E(x))} ^2""=(3^2*{1 \over 6}) +(6^2*{1\over 2})+(9^2*{1 \over 3})- 6.5^2"

"var(x) =4.25"


"var(y) = 9(4.25) = 38.25"

b)

"E(W) = 2E(X)-4E(Y)+5""E(X) = np = 6({1 \over 3}) =2""E(Y) ={1 \over 2}"

"E(W) =(2*2)- 4({1\over 2})+5"

E(W) = 7


"var(w) = 2^2 var(x) + 4^2 var(y) - 2(2*4) cov(xy)"

But "Cov(xy) =0"


"\therefore var(w) =4var(x)+16var(y)"

"Var(x) =np(1-p) =6*{1 \over 3}*{2 \over 3}={4 \over 3}"

"Var(y) ={1\over 4}"


"Var(W) =4({4 \over 3}) +16({1 \over 4})"

Var(W) =9.33333333